Car A and car B both travel 1320 miles. Car A drove at an average speed of 10 miles per hour faster than car B and made the trip in one hour less time. Find the average speeds of both cars.
Answer by richwmiller(17219)

r*t=d
r*t=1320
(r+10)*(t-1)=1320
r=110 and t=12
r+10=120 t-1=11